adovbush:

Given the function f(x)= 3+|x-2|, find the turning point. What x value makes the absolute value equal to 0? Then, substitute that value in to find the ordered pair. and Given the function f(x)= 3+|x-2|, find f(-4).

adovbush:

@Falconmaster @dude @Shadow

dude:

` find the turning point. *`? What is *

adovbush:

Oh that's not needed sorry

dude:

https://www.desmos.com/calculator/2qcbp28wfx Turning point in this case is just the vertex

dude:

For the second part x-2=0 Solve for x

adovbush:

2?

dude:

Right so that is when the absolute value is 0 tranq you can input too if you think theres a better way to say it or to check

adovbush:

I think I got it.... The turning point is (3,3). The x to make the absolute value negative would be 2. The ordered pair would then be (0,3)

adovbush:

But I dunno the second one.

dude:

Check your turning point, refer to the graph

adovbush:

Idk dude

dude:

-mk3iUmBQsuR5ZOv4esKXQ.png

adovbush:

That's the answer?

dude:

lol, yes that is the turning point

adovbush:

Thanks dude, you da best x'D

dude:

We didnt finish part 2

dude:

`Given the function f(x)= 3+|x-2|, find f(-4).`

dude:

f(-4) just means y when x= -4 \(f(-4)= 3+|(-4)-2|\) \(f(-4)= 3+|(-6|\) \(f(-4)= 3+6\) \(f(-4)=\) answer

adovbush:

"Thanks dude, you da best" x'D

dude:

Cools :P